40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 38 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  15194   Thu Feb 6 21:54:13 2020 JonUpdatePSLc1psl bench testing complete

Today I engineered the last piece of the new c1psl system: the multi-bit binary output (mbbo) channels that control the MC servo board gains. These 6-bit channels have to be split across two 4-bit Acromag registers. To enforce synchronous switching, I adapted the latch.py script developed by Gautam to address this problem in c1iscaux. Analogously to the c1iscaux implementation, I scripted the code to automatically run as a systemd service which is launched by the main modbusIOC service. I tested this all using the DB37 LED test board and confirmed it to work.

This now completes the electronics bench testing.

There are still several DB37 connectors to be wired, which carry only spare channels for future use and are not interfaced with the EPICS IOC. Jordan and I discussed this today and he or Chub will complete it shortly. To allow time for the spare channel wiring to be completed (as well as for more locking progress before interruption), Gautam and I think Monday/Tuesday next week would be the earliest possible window to install the new system.

  15193   Thu Feb 6 16:14:44 2020 ranaUpdateGeneraloffice area temperature

I changed the office area thermostate near Steve's desk from 68F to 73F today. Please do not change it.

If anyone from facilities comes to adjust something, please put the details in the elog on the same day so that we can know to undo that change rather than chase down other drifts in the system.

  15192   Thu Feb 6 01:25:50 2020 gautamUpdateLSCLocking updates

Summary:

I managed to partially stabilize the arm citculating powers - they stay in a region in which the REFL 11 signal is hopefully approximately linear and so I can now measure some loop TFs and tweak the transition appropriately. 

Procedure:

The main change I made tonight was to look at the REFL11 signal as I swept the ALS CARM offset through 0. I found that the maximum arm powers coincided with a non-zero REFL11 signal value (i.e. a small CARM offset was required at the input to the CARM_B filter bank). Not so long ago, I had measured the PM/AM ratio for 11 MHz to be ~10^5 - so it's not entirely clear to me where this offset is coming from. Then, I was able to turn on the integrator (z:p = 20:0) in the CARM_B filter bank while maintaining high POP_DC. At this point, I ramped up the IN2 gain on the IMC servo board (= AO path), and was able to further stabilize the power. 

Attachment #shows this sequence from earlier in the evening. Note that in this state, both ALS and IR control of CARM is in effect. The circulating power is fluctuating wildly - partly this is probably the noisy ALS control path, but there is also the issue of the (lack of) angular control - although looking at the transmon QPDs and the POP QPD signals, they seem pretty stable.

The next step will be to try and turn off the ALS control path. Eventually, I hope to transition DARM control to AS55 as well. But at this point, I can at least begin to make sense of some of the time series signals, and get some insight into how to improve the lock.

Quote:

No systematic diagnosis scheme comes to mind.

Attachment 1: semiStableArms.png
semiStableArms.png
Attachment 2: armAngStability.png
armAngStability.png
  15191   Thu Feb 6 01:16:58 2020 gautamUpdateLSCDiagnosis results

Summary:

I did some more detailed tests to see if I could isolate where the excess ALS noise at low CARM offset is coming from, by measuring the spectrum of the IMC error point (in loop). The results, shown in Attachment #1 and #2, are inconclusive.

Details:

Since MC_F didn't show any signatures of elevated noise, I decided to hook up an SR785 to the A excitation bank TEST1 input of the IMC servo board to monitor the in-loop error signal. I initially took a few measurements spanning 800 Hz in frequency, and to my surprise, I found that there was elevated noise in the frequency band we see an increase in the ALS noise, even when the CARM feedback goes to the ETMs (so the IMC cavity is in principle isolated from the main interferometer). This is Attachment #1. So I re-took a couple of measurements (this time only for the case of CARM feedback to the ETMs), with a 200 Hz frequency span, and found no significant noise elevation. This is Attachment #2. I am led to conclude that the IMC error point level changes over time for reasons other than the CARM offset - it'd be nice to have a spectrogram of the IMC error point and compare excursions relative to the median level over a few 10s of minutes, but we don't have this data stream digitized by the CDS system - maybe I will hijack the MC_L channel temporarily to record this data stream. It seems a waste that we're not able to take full advantage of the measured <10pm RMS noise of the IR ALS system.

Attachment 1: IMCspec_ALS.pdf
IMCspec_ALS.pdf
Attachment 2: IMCspec_ALS_smallSpan.pdf
IMCspec_ALS_smallSpan.pdf
  15190   Wed Feb 5 21:13:17 2020 YehonathanUpdateIOOIMC Ringdowns extended data analysis

I translate the results obtained in the previous elog to the IMC 3 mirror cavity. I assume the loss in each mirror in the IMC is equal and that M2 has a negligible transmission.

I find that to a very good approximation the loss per IMC mirror is 2/3 the loss per mirror in the 2 mirror cavity model. That is the loss per mirror in the IMC is 56 ppm. The transmission per mirror in the IMC is the same as in the 2 mirror model, which is 1980 ppm.

The total transmission is the same as in the 2 mirror model and is given by:

\frac{P_0}{P_0+P1}KT^2\approx 90\%

where \frac{P_0}{P_0+P1} is the coupling efficiency to the TEM00 mode.

  15189   Wed Feb 5 21:04:10 2020 YehonathanUpdatePSLBench testing of PSL ai channels

{Yehonathan, Jon}

We retested the failed ai channels. Most of them got fixed by applying the inverse calibration in the yaml file.

We still find some anomalous channels, mostly in the DB25 connector. Turns out, their limits were ill-defined in the EPICS database. Specifying the right limit fixed the issue.

We find one miswired channel (BNC4). We connected the BNC to the right channel on the Acromag unit which fixed the issue.

Overall all the ai channels were successfully bench-tested.

Quote:

I checked the failed channels against the EPICS database definitions and the yaml file inputted to IFOTest. The channels where the readings are something other than +10/0 V, but the high/low values do change, I think can be attributed to one of two things:

  • An incorrect gain and/or offset conversion parameter in the yaml file
  • The EPICS SMOO parameter (smoothing) is set to some long value

I fixed the channel gains/offsets in the master yaml file (PSL.yaml). I also disabled smoothing in the EPICS defintions of the new PSL channels for the purpose of testing. We can uncomment those lines after installing the new chassis if noise is a problem. Please go ahead and re-test the channels again.

Quote:

I marked in red the failed channels. Some of them are probably calibration issues, but the ones that show the same voltage for high and low are probably disconnected wires.

  15188   Wed Feb 5 16:35:12 2020 gautamUpdateLSCDiagnosis plan

The goal is to try and identify the source of the excess ALS noise as the CARM offset is reduced. The idea is to look at the MC_F spectrum (or the IMC error point) in a few conditions:

  1. Regular CARM --> MC2 actuation scheme, PRMI locked on 3f signals, CARM held off resonance.
  2. Regular CARM --> MC2 actuation scheme, PRMI locked on 3f signals, CARM held on resonance.
  3. Alternate CARM --> 7.5*ETMX + 1.5*ETMY, PRMI locked on 3f signals, CARM held on resonance.
  4. Control arms in X/Y basis, lock PRMI on 3f signals and bring the arms into resonance individually, look for excess ALS noise.

#1 vs #2 is like a control experiment, we expect to see the excess noise imprinted on the MC length and hence in MC_F (provided the sensing noise is low enough). #2 vs #3 will be informative of something like backscatter to the PSL increasing the frequency noise. #2/3 vs #4 will help isolate the problem to an individual arm's AUX PDH loop or some optomechanical effect.

I was looking back at some spectra from the last couple of nights but I don't really have an apple-to-apple comparison in the various actuation schemes (some ALS loops were engaged/disengaged), so I'll do a more systematic test tonight. Already, it looks like MC_F is not a good candidate to look for the excess frequency noise, I don't really see a big difference between conditions #1 and #2. According to this, we are looking for an increase at the level of a few 100Hz/rtHz @ ~40 Hz, wheras MC_F is much noisier.

Attachment 1: ALSnoise.pdf
ALSnoise.pdf
  15187   Wed Feb 5 08:57:11 2020 YehonathanUpdatePSLBench testing of PSL ai channels

I checked the failed channels against the EPICS database definitions and the yaml file inputted to IFOTest. The channels where the readings are something other than +10/0 V, but the high/low values do change, I think can be attributed to one of two things:

  • An incorrect gain and/or offset conversion parameter in the yaml file
  • The EPICS SMOO parameter (smoothing) is set to some long value

I fixed the channel gains/offsets in the master yaml file (PSL.yaml). I also disabled smoothing in the EPICS defintions of the new PSL channels for the purpose of testing. We can uncomment those lines after installing the new chassis if noise is a problem. Please go ahead and re-test the channels again.

Quote:

I marked in red the failed channels. Some of them are probably calibration issues, but the ones that show the same voltage for high and low are probably disconnected wires.

  15186   Tue Feb 4 18:13:01 2020 YehonathanUpdatePSLBench testing of PSL ai channels

{Yehonathan, Jon, Jordan}

I tested the ai channels of the new PSL Acromag by looping an already-tested ao channel (C2:PSL-FSS-INOFFSET) back to the different ai channels.

I use Jon's IFOTest with /users/jon/ifotest/PSL.yaml.

I created a spreadsheet for the testing based on the current wiring spreadsheet. I added two columns for the high and low readings for each ai channel (attached pdf).

I marked in red the failed channels. Some of them are probably calibration issues, but the ones that show the same voltage for high and low are probably disconnected wires.

I redid the test on the channel that seemed disconnected to confirm.

I created a yaml file with all the failed channels for retesting called /users/jon/ifotest/PSL_failed_channels.yaml.

Attachment 1: c1psl_wire_testing_-_By_Connector.pdf
c1psl_wire_testing_-_By_Connector.pdf c1psl_wire_testing_-_By_Connector.pdf c1psl_wire_testing_-_By_Connector.pdf c1psl_wire_testing_-_By_Connector.pdf
  15185   Tue Feb 4 02:13:02 2020 gautamUpdateLSCLocking updates

Summary:

The CARM-->RF transition remains out of reach. No systematic diagnosis scheme comes to mind.

Details:

  • Config is PRFPMI, SRM is misaligned macroscopically.
  • PRMI can easily be locked with 3f signals while CARM is offset from resonance. Aided by DAFI, I turned on the PR violin filter in the BS output section to prevent it from ringing up, making the lock much more robust.
  • When the CARM offset is reduced
    • POP22 level dips and sometimes goes negative - i don't see this in my simple simulations. POP22 is the trigger signal for MICH/PRCL loops, so to prevent the PRMI lockloss, I mix in some POPDC into the trigger matrix element.
    • Once the circulating power exceeds ~10, the ALS noise apparently increases.
    • The arms "buzz" through resonance, but the power fluctuation is nearly 0-200 in TRX/TRY, corresponding to several CARM linewidths, but all the out-of-loop ALS noise measurements have me believe that we are close to the CARM linewidth in noise. So we should only see ~factor of 2 fluctuation in power.
    • The RF error signal for CARM (=REFL 11) doesn't show any features that i can use to aid the transition / diagnose what is going on systematically.
  • Koji suggested changing the actuation for CARM from MC2 to the ETMs, and check if the MC OSEMs witness the excess motion at small CARM offsets
    • The ALS transition is scripted, so I had to make a modified version that accommodates this changed actuation scheme.
    • The usual CARM-->MC2 matrix element is -1. 
    • The frequency actuation strength of MC2 is ~3x that of the ETMs. Additionally, ETMX has 5x the series resistance of ETMY. So I used the output matrix elements shown in Attachment #1 so as to get the same loop UGF with the same loop gains elsewhere in the chain. Confirmed the actuation strength is the same using the sensing matrix infrastructure and comparing line heights.
    • Attachment #2 shows the measured UGF - both CARM and DARM look okay to me.
  • With this new ALS output matrix actuation scheme, I was able to make it to PRMI + arms on zero offset a couple of times tonight, but the drifting input alignment makes the PRMI lock not so robust anymore.

TBC. Mercifully at least the shaker stayed still tonight.

Attachment 1: modifiedOutMat.png
modifiedOutMat.png
Attachment 2: OLTFs.pdf
OLTFs.pdf
  15184   Mon Feb 3 15:22:39 2020 JonUpdatePSLc1psl progress/Acromag ADC grounding

I tested the c1psl AO channels on the electronics bench on Friday. While I found all the wiring to be correct, some of the channels exhibited excess noise with all appearances of a grounding problem.

Today Jordan, Gautam, and I investigated this further. It is indeed a grounding problem, but actually with the Acromag ADCs. The Acromag DAC outputs are single-ended (return is grounded), so (for the purpose of a loopback test) I would expect to leave the ADC inputs ungrounded. This is the configuration I tested Friday. Today we also tested driving the ADC with a floating source. The ADC noise behavior is exactly the same, whether the source end is grounded or not.

However, grounding the minus pin of the ADC channel eliminates the noise. We don't understand why this seems to be required irrespective of the driving source, so there something we're missing about the ADC design. As it turns out, this same fix was made to the AI channels of the previously-upgraded Acromag machines. I know Chub and I had to do this for the AI channels of c1vac, but at the time we thought the source grounding was causing the issue. However, today Jordan and I looked inside c1iscaux, which Chub wired, and confirmed that its AI channels are wired in the same way.

So in any case, Jordan is grounding the c1psl AI channels in the same way as c1iscaux. Once this is done, we'll continue with the bench testing tomorrow.

gautam: here are my notes about this issue when i was doing the c1iscaux testing. As I note there, "previously-upgraded Acromag machines" in the plural may be a bit of a stretch - I have no idea what the grounding situation is in c1susaux / c1auxex for example.

  15183   Mon Feb 3 13:54:10 2020 YehonathanUpdateIOOIMC Ringdowns extended data analysis

I extended the ringdown data analysis to the reflected beam following Isogai et al.

The idea is that measuring the cavity's reflected light one can use known relationships to extract the transmission of the cavity mirrors and not only the finesse.

The finesse calculated from the transmission ringdown shown in the previous elog is 1520 according to the Zucker model, 1680 according to the first exponential and 1728 according to the second exponential.

Attachment 1 shows the measured reflected light during an IMC ringdown in and out of resonance and the values that are read off it to compute the transmission.

The equations for m1 and m3 are the same as in Isogai's paper because they describe a steady-state that doesn't care about the extinction ratio of the light.

The equation for m2, however, is modified due to the finite extinction present in our zeroth-order ringdown.

Modelling the IMC as a critically coupled 2 mirror cavity one can verify that:

m_2=P_0KR\left[T-\alpha\left(1-R\right)\right]^2+\alpha^2 P_1

Where P_0 is the coupled light power 

P_1 is the power rejected from the cavity (higher-order modes, sidebands)

K=\left(\mathcal{F} /\pi \right )^2 is the cavity gain.

R and T are the power reflectivity and transmissivity per mirror, respectively.

\alpha^2 is the power attenuation factor. For perfect extinction, this is 0.

Solving the equations (m1 and m3 + modified m2), using Zucker model's finesse, gives the following information:

Loss per mirror = 84.99 ppm
Transmission per mirror = 1980.77 ppm
Coupling efficiency (to TEM00) = 97.94%
Attachment 1: IMCTransReflAnalysis_anotated.pdf
IMCTransReflAnalysis_anotated.pdf
  15182   Fri Jan 31 16:57:09 2020 gautamUpdateGeneralMetal PMC parts

Jon brought over a box of parts for constructing the metal PMCs. I have stored it along the Y-arm, on top of the green optics cabinet.

I didn't do an exhaustive inventory check, but the following are the rough contents of the box:

  • 41 deg AoI flat mirrors, R=99% @ 1064nm --- 11 pcs
  • 6.8 deg AoI curved mirrors --- 5 pcs
  • PZTs --- 3pcs
  • Metal PMC body --- 2 pcs
  • "Baked PZT endcaps" --- 3 pcs
  • Ball bearings, clamps, misc hardware

I didn't inspect the optics but since we have so many, I am hoping we can find 3 good quality ones for one cavity at least. We should check that the geometry is suitable for our RF sideband frequencies.  

  15181   Fri Jan 31 16:04:30 2020 gautamUpdateIOOInput pointing drift

One factor which hampers locking efforts is the apparent drift of the input beam into the IFO. Over timescales of ~1 hour, I have noticed that the spot on the AS camera drifts significantly (~1 spot size) in pitch. The IPPOS QPD bears out this observation, see Attachment #1. The IMC WFS control signals do not show a correlated drift, hence my claim that the TTs are to blame.

I am able to correct this misalignment by moving TT1 in pitch (see Attachment #2, which shows some signals from a ~1 hour PRMI lock, during which time the pointing drifted, and I corrected it by moving TT1 pitch). Assuming the problem is purely TT1 pitch drifting, this corresponds to 3mm / 6m ~500urad of shift in 1 hour - seems very large. The fact that the drift is only present in pitch and doesn't really show up in yaw makes me think the problem is likely mechanical (unless the voltage to the top two coils is drifting relative to the bottom, but no LR drift, which would be very coincidental). At the moment, this is just an annoyance, but it'd be good for this problem to be fixed.

I haven't yet figured out how to make ndscope export these plots to SVG preserving the dark color theme, hence the weird light axes...

Attachment 1: IPdrift.pdf
IPdrift.pdf
Attachment 2: IPdrift_PRMI.pdf
IPdrift_PRMI.pdf
  15180   Thu Jan 30 22:02:42 2020 shrutiUpdateGeneraldelay line frequency discriminator for PM

I could not find any level 3 mixers, but by adjusting the beat frequency the power in each of the delay line channels rose to almost 6.5 dBm.

In short: Delay line seems to work

Things I did earlier today:

  1. Played with the slow servo on the FSS screen, but then reset the parameters to what was there before (Later found out that this was to lock the PSL freq to the IMC when the IMC power is significant.)
  2. Connected the AG 4395A to the X PZT
  3. Closed the PSL shutter

Transfer function measurement: (Refer Attachment 1)

Everything about the setup remained as I had left it earlier: described in elog 15174

except

  • SR560 gain set to 10, DC coupled
  • DC block at channel A of Agilent (The measurement was A/R)

I did not use a slow servo, but took individual sweeps adjusting the PSL temperature each time to bring the error voltage between +/-25 mV. The beat frequency was over 100 MHz.

For the plot posted in Attachment 1, the measurement paramters are the following. Will do further measurements/analysis tomorrow.

# AG4395A Measurement - Timestamp: Jan 30 2020 - 21:58:00
# Parameter File: TFAG4395Atemplate.yml
#---------- Measurement Parameters ------------
# Start Frequency (Hz): 50000.0, 50000.0
# Stop Frequency (Hz): 1000000.0, 1000000.0
# Frequency Points: 801, 801
# Measurement Format: LOGM, PHAS
# Measuremed Input: AR, AR
#---------- Analyzer Settings ----------
# Number of Averages: 1
# Auto Bandwidth: Off, Off
# IF Bandwidth: 1000.0, 1000.0
# Input Attenuators (R,A,B): 0dB 0dB 0dB 
# Excitation amplitude = -20.0dBm

Quote:

yes, its fine to use this with a level 3 or level 7 mixer; let's see some PM transfer functions !

Quote:

Is this sufficient enough for the mixer to work?

Attachment 1: Figure_2.png
Figure_2.png
  15179   Thu Jan 30 17:41:10 2020 gautamUpdatePSLErrant FSS_INOFFSET change

You can trend the data for the past few hours and see what the appropriate value. I think these tests should only be done when whoever is running a test is in the lab.

P.S. I was surprised that the IMC didn't lose lock when this step was applied. I manually stepped this voltage between +/- 10 V and didn't see any response in the FSS readbacks. Either the channel doesn't work, or there is a divide by 40 in the physical circuit or something...

Quote:

A script I was testing errantly set C1:PSL-FSS_INOFFSET => 10 V at about 5:30 pm. I manually reverted the channel value to 0, but I don't know what the value was initially. Someone please check this value if there are problems locking the FSS.

  15178   Thu Jan 30 17:31:28 2020 JonUpdatePSLErrant FSS_INOFFSET change

A script I was testing errantly set C1:PSL-FSS_INOFFSET => 10 V at about 5:30 pm. I manually reverted the channel value to 0, but I don't know what the value was initially. Someone please check this value if there are problems locking the FSS.

  15177   Thu Jan 30 15:24:10 2020 ?UpdateGeneraldelay line frequency discriminator for PM

yes, its fine to use this with a level 3 or level 7 mixer; let's see some PM transfer functions !

Quote:

Is this sufficient enough for the mixer to work?

  15176   Thu Jan 30 12:52:10 2020 JonUpdateBHDMetal OMCs procured

Last night Yehonathan and I located the two steel PMCs in the QIL, with help from Anchal. They are currently sitting on my desk in Bridge, inside a box that also contains optics and other OMC parts. I will bring them over to the 40m the next time I come.

  15175   Wed Jan 29 12:40:24 2020 YehonathanUpdateIOOIMC Ringdowns preliminary data analysis

I analyze the IMC ringdown data from last night. 

Attachment 1 shows the normalized raw data. Oscillations come in much later than in Gautam's measurement. Probably because the IMC stays locked.

Attachment 2 shows fits of the transmitted PD to unconstrained double exponential and the Zucker model.

Zucker model gives time constant of 21.6us

Unconstrained exponentials give time constants of 23.99us and 46.7us which is nice because it converges close to the Zucker model.

Attachment 1: IMCRingdownNormalizedRawdata.pdf
IMCRingdownNormalizedRawdata.pdf
Attachment 2: IMCTransPDFits.pdf
IMCTransPDFits.pdf
  15174   Wed Jan 29 12:29:33 2020 shrutiUpdateGeneraldelay line frequency discriminator for PM

 Today I began working on a TF measurement based on the delay line frequency discriminator setup in elog 4254 using a single mixer (without the 'I' and 'Q' readout).

For this, I re-organised the setup for the PLL measurement of the transfer function (elog 15148), increasing the HEPA for the initial changes while the PSL door was open, and then reverting it back to ~30%:

  • I removed the 20dB coupler and connected the splitter directly after the amplifer to split the beat note signal into two coaxial cables one of which was ~1.5m longer than the other
  • The recombined signals were combined in a mixer outside the PSL enclosure. I also replaced the 1.9 MHz LPF with a 5 MHz LPF.
  • I used an SR 560 to amplify the signal after the LPF.

With the above setup the power that was seen at each channel of the delay line was <1dBm, which is not ideal for the any of the available mixers.

After the group meeting, I changed the amplifer to ZHL-3A (that is near the beat mouth) instead of a ZFL-500HLN because it had a higher gain (~28dB as opposed to ~19dB of the latter). The power seen at each of the delay line channels is over 5.5 dBm. This is consistent with the estimation 0 dBm beat -> -20 dBm after 20dB coupler -> 8 dBm after amplifier -> 5 dBm after splitter with insertion loss of 3 dB.

Is this sufficient enough for the mixer to work? In Attachment 1: A shows the mixer output (point B in Attachment 2) when the IMC is locked, in B the IMC is unlocked at the middle of the spectrum, and each of the dips show the DC voltage being sent to the PSL temperature servo being decreased by 0.01 V.

Gautam pointed me to the location of a few other RF amplifiers (ZHL-32A+, ZHL-1A) which don't possess a higher gain but can be used without disrupting the ALS related work (I was told).

For shorter duration changes that I made later, I opened and closed the PSL enclosure doors without changing the HEPA.

Attachment 2 shows the current setup as is, but I might add a PSL servo tomorrow to stabilise its frequency corresponding to a null mixer output without driving anything else.

Attachment 1: 20200128.png
20200128.png
Attachment 2: IMG_BB01C068495A-1.jpeg
IMG_BB01C068495A-1.jpeg
  15173   Wed Jan 29 03:05:47 2020 rana, gautamUpdateSUSMC misalignments / sat box games

In the last couple days, as the IMC ringdowns have been going on, we have noticed that the MC is behaving bad. Misaligning, drifting, etc.

Gautam told me a horror story about him, Koji, and melted wires inside the sat boxes.

I said, "Its getting too hot in there. So let's take the lids off!"

So then we:

  1. Removed the lid (only 4 screws were still there)
  2. cut off some of the shield - ground wires and insulated them with electrical tape
  3. squished the IDC connectors on tightly
  4. left it this way to see if MC would get better - certainly the painfully hot heatinks inside the box were now just 110 F or so

After some minutes, we saw no drifting. So maybe my theory of "hot heatsink partially shorting a coil current to GND through partially melted ribbon cable" makes sense? IF this seems better after a month, lets de-lid all the optics.

Let's look at some longer trends and be very careful next to MC2 for the next 3 days! I have put a dangerous mousetrap there to catch anyone who walks near the vacuum chamber.

gautam: the grounding situation per my assessment is that the shield of all the IDC cables are connected to a common metal strip at 1X5 - but in my survey, I didn't see any grounding of this strip to a common ground.

Attachment 1: IMG_8366.JPG
IMG_8366.JPG
  15172   Wed Jan 29 00:29:43 2020 gautamUpdateLSClocking 2020

The goal tonight was to go through the locking scripts to see if I could recover the state from November 2019, when I could have the arm lengths controlled by ALS, and sit at zero CARM offset with the PRMI remaining locked and the arm powers fluctuating between 0-300. The IR-->ALS transitions went smoothly tonight, and the PRMI locking was also fairly robust when the CARM offset was large, but was less good when reduced to 0. Nevertheless, it is good to know that the system can be restored to the state it was late last year. Next step is to figure out how to keep the PRMI locked and get the AO path engaged, this was what I was struggling with before the new year.

Attachment 1: PRFPMI_2020Jan.png
PRFPMI_2020Jan.png
  15171   Wed Jan 29 00:27:13 2020 gautamUpdateComputer Scripts / Programsmcup / mcdown modified

To fix the apparent slowness of execution of the caput commands on megatron, I changed the "ewrite" macro in the mcup and mcdown scripts to use ezcawrite instead of caput. The old lines are simply commented out, and can be reverted to at any point if we so desire. After these changes, we saw that both scripts complete execution much faster.

  15170   Tue Jan 28 20:51:37 2020 YehonathanUpdateIOOIMC WFS servos stable again

I resume my IMC ringdown activities now that the IMC is aligned again.

To avoid any accidental misalignments Gautam turned off all the inputs to the WFS servo.

I set up a PD and a lens as in attachment 1 (following Gautam's setup).

I connect the REFL, TRANS and INPut PDs to the oscilloscope.

I connect a Siglent function generator to the AOM driver. I try to shut off the light to the IMC using 1V DC waveform and pressing the output button manually. However, it produced heavily distorted step function in the PMC trans PD.

I use a square wave with a frequency of 20mHz instead with an amplitude of 0.5V offset of 0.25V and dutycycle of 1% so there will be minimal wasted time in the off state. I get nice ringdowns (attachment 2) - forgot to take pictures. The autolocker slightly misaligns the M2 every time it is acting, so I manually align it everytime the IMC gets unlocked.

Data analysis will come later.

I remove the PD and lens and reenable the WFS servo inputs. The IMC locks easily. The WFS outputs are very different than 0 now though.

  15169   Tue Jan 28 19:40:15 2020 shrutiUpdateGeneralPLL / PM measurement of Xend NPRO PZT

Over the past few days, I have been trying to make measurements of the phase modulation transfer function by modulating the X end laser PZT via PLL.

The setup was modified every time during the experiment in the same manner as mentioned in elog 15148.

I could not make the PLL lock for long enough to take a proper TF measurement, resulting in TFs that look like Attachment 1. The next step would be to use the method of a delay line frequency discriminator instead of the PLL.


Comments about locking with LB1005 PI controller:

  1. I do not understand why the high PI corner frequency of 1kHz or 3kHz was required to lock.
  2. The rms level of the error signal when locked was ~100 mV, which is 25% of the total mixer range (~400 mVpp). Decreasing the gain only caused the loop to go out of lock and did not decrease this noise in the error signal.
  3. The setup was also partly inside the PSL enclosure, with the HEPA turned to 100%, which is probably a noisy environment for this measurement. Closing and opening the shutters or any disturbance near the enclosure resulted in movement of the beat note up to 5 MHz.
  4. It may have been a better idea to actuate the PSL laser instead of the X NPRO because of its larger range, but would this solve the issue with the noise?
Attachment 1: PMTF.pdf
PMTF.pdf
Attachment 2: BeatSpectrum.pdf
BeatSpectrum.pdf
  15168   Tue Jan 28 19:12:30 2020 JonConfigurationPSLSpare channels added to c1psl chassis

After some discussion with Gautam, I decided to build more spare channels into the new c1psl machine. This is anticipation of adding new laser and ISS channels in the near future, to avoid having to disconnect the installed chassis and pull it out of the rack. The spare channels will be wired to DB37M feedthroughs on the front side of the chassis, with enough wire length to be able to pull the breakout boards out of the front to reconfigure their wiring as needed (e.g., split off channels onto a separate connector).

To have enough overhead, this will require installing 1 additional ADC unit (XT1221) and 1 additional DAC (XT1541). We have enough spare BIO channels among the existing units (both sinking and sourcing). This will give us:

  • 13 spare ADC channels
  • 14 spare DAC channels
  • 16 spare sinking BIO channels
  • 12 spare sourcing BIO channels

The updated c1psl chassis wiring assignments are attached. It adds 4 new DB37M connectors for the spare channels (highlighted in yellow) and fixes one typo Jordan found while wiring today. The most current spreadsheet is available here.

Attachment 1: c1psl_feedthrough_wiring_v2.pdf
c1psl_feedthrough_wiring_v2.pdf c1psl_feedthrough_wiring_v2.pdf c1psl_feedthrough_wiring_v2.pdf c1psl_feedthrough_wiring_v2.pdf c1psl_feedthrough_wiring_v2.pdf
  15167   Tue Jan 28 17:36:45 2020 gautamConfigurationComputersLocal EPICS7.0 installed on megatron

[Jon, gautam]

We found that the caput commands were taking much longer to execute on megatron than on pianosa (for example). Suspecting that this had something to do with the fact that megatron was using EPICS binaries from the shared NFS drive which were compiled for a much older OS, I installed the latest stable release of EPICS on megatron. The new caput commands execute much faster. I also added the local EPICS directory to the head of the $PATH variable used by the MC autolocker and FSS Slow scripts, so that they use the new caput command. But mcup is still slow - maybe my new path definition isn't picked up and it is still using the NFS binaries? To be looked into...

Quote:

There were a bunch of medm processes stalled on megatron (connected with screenshot taking). To see if they were interfering with the other scripts, I killed all of the medm processes, and commented out the line in the crontab that runs the screenshots every 10 mins. Let's see if this improves stability.

  15165   Tue Jan 28 16:01:17 2020 gautamUpdateIOOIMC WFS servos stable again

With all of the shaking (man-made and divine), it was a hard to debug this problem. Summary of fixes:

  1. The beam was misaligned on the WFS 1 and 2 heads, as well as the MC2 trans QPD. I re-aligned the former with the IMC unlocked, the latter (see Attachment) with the IMC locked (but the MC2 spot centering loops disabled).
  2. I reset the WFS DC and RF offsets, as well as the QPD offsets (once I had hand-aligned the IMC mirrors to obtain good transmission).

At least the DC indicators are telling me that the IMC locking is back to a somewhat stable state. I have not yet checked the frequency noise / RIN.

Attachment 1: QPD_recenter.png
QPD_recenter.png
  15164   Tue Jan 28 15:39:04 2020 gautamConfigurationComputersSluggish megatron?

There were a bunch of medm processes stalled on megatron (connected with screenshot taking). To see if they were interfering with the other scripts, I killed all of the medm processes, and commented out the line in the crontab that runs the screenshots every 10 mins. Let's see if this improves stability.

  15163   Tue Jan 28 14:33:24 2020 gautamUpdatePSLInferred free-running frequency noise

To conclude my PMC noise investigations: Attachment #1 shows the PMC noise inferred from the calibrations earlier in this thread and the fitted OLTF for the PMC loop. Attachment #2 compares the frequency noise (inferred from the error point of the PMC servo) when the IMC is locked / unlocked. I don't know what to make of the fact that the PMC suggests improvement from ~20 Hz onwards already - does this mean that the NPRO noise model is wrong by 1 order of magnitude at 30 Hz?

  • The IMC was locked for the measurement shown in Attachment #1.
  • The in-loop spectra of the error (at the I/F output of the mixer) and control (at TP3) signals were measured with the SR785.
  • The control signal voltage monitors don't seem to work - neither the front panel LEMO nor the signals hooked up to the CDS system show me sensible shapes for the spectra between 1-3 Hz.
  • To convert in loop to free-running, I multiplied the measured error (control) signal spectra by \left | 1-L \right | (\left | \frac{L}{1-L} \right |), where L is the OLTF. THe control signal was pre-processed by multiplying by a pole at 11.3 Hz, corresponding to the LPF formed by the 63.3 kohm series resistor and the 225 nF PZT capacitance.
  • The "NPRO noise model" curve is 10^4/f Hz/rtHz.

While I initially thought the 1/f^2 rise below ~100 Hz is attributable to the IMC cavity length fluctuations, I found that this profile is present even in the measurement with the PSL shutter closed. I am not embarking on a detailed PMC noise budgeting project for now. Note however that we are not shot noise limited anywhere in this measurement band.

  • The measured TF (dots in Attachment #5) was fit with LISO (solid lines in Attachment #5) to allow inferring the out-of-loop servo noise by monitoring the in-loop noise (that plot to follow).
Attachment 1: inLoopNoise_IMClocked.pdf
inLoopNoise_IMClocked.pdf
Attachment 2: freqNoiseComparison.pdf
freqNoiseComparison.pdf
  15162   Tue Jan 28 08:26:53 2020 ranaFrogsPEMshaking

https://breakthrough.caltech.edu/magazine/2019-aug/#article-Listening-with-Light

  15161   Mon Jan 27 21:48:49 2020 gautamUpdatePSLRingdown measurements

It's fine to block the WFS while doing ringdowns but please return the config to normal so I don't have to spend time every night recovering the interferometer before doing the locking. As I mention in that post, it is possible to do this in a non-invasive way without having to run any extra cables / permanently block any beams. If there is some issue with the data quality, then we can consider a new setup. But I see no reason to re-invent the wheel.

The IMC was also massively misaligned. I had to re-align both MC1 and MC2 to recover the lock. I took this opportunity to reset the WFS offsets. Please do not disturb the alignment of the existing optical layout unless you verify that everything is working as it should be after your changes.

And for whatever reason, ITMX was misaligned. If you do something with the interferometer, no matter how minor it seems, please leave a note on the ELOG. It will save many painful debugging hours.

As I fix these, the seismic activity has gone up no. I'll wait around for an hour, but not an encouraging restart to the locking 😢 

Quote:

Zeroth order IMC ringdown

Following Gautam's IMC ringdown setup, I took the the REFL PD form the PMC ringdown experiment and installed it in the IMC REFL path blocking WFS2 (Attachment 1).

Attachment 1: elevatedSeis.pdf
elevatedSeis.pdf
  15160   Mon Jan 27 21:35:06 2020 YehonathanUpdatePSLRingdown measurements

Zeroth order IMC ringdown setup

Following Gautam's IMC ringdown setup, I took the REFL PD form the PMC ringdown experiment and installed it in the IMC REFL path blocking WFS2 (Attachment 1).

I also ran a BNC cable from the transmission PD that Gautam installed on the IMC table to the vertex where the signals are measured on the scope. 

I offloaded the WFS servo output values onto the MC alignment (using the WFS servo relief script) so that its dc values would be correct when the servo is off.

Unfortunately, it seems like the script severely misaligned the MC mirrors at some point when the MC got unlocked. We should fix the script such that it stops when the offloading is complete.

We got the MC realigned but left it in a state where it is not locking easily.

Attachment 1: IMC_REFL_Beam_Path.jpg
IMC_REFL_Beam_Path.jpg
  15159   Mon Jan 27 18:16:30 2020 gautamConfigurationComputersSluggish megatron?

I've also been noticing that the IMC Autolocker scripts are running rather sluggishly on Megatron recently. Some evidence - on Feb 11 2019, the time between the mcup script starting and finishing is ~10 seconds (I don't post the raw log output here to keep the elog short). However, post upgrade, the mean time is more like ~45-50 seconds. Rana mentioned he didn't install any of the modern LIGO software tools post upgrade, so maybe we are using some ancient EPICS binaries. I suspect the cron job for the burt snapshot is also just timing out due to the high latency in channel access. Rana is doing the software install on the new rossa, and once he verifies things are working, we will try implementing the same solution on megatron. The machine is an old Sun Microsystems one, but the system diagnostics don't signal any CPU timeouts or memory overflows, so I'm thinking the problem is software related...

Quote:

The burt snapshotting is still not so reliable - for whatever reason, the number of snapshot files that actually get written looks random. For example, the 14:19 backup today got all the snaps, but 15:19 did not. There are no obvious red flags in either the cron job logs or the autoburt log files. I also don't see any clues when I run the script in a shell. It'll be good if someone can take a look at this.

  15158   Mon Jan 27 14:01:01 2020 JordanConfigurationGeneralRepurposed Sorenson Power Supply

The 24 V Sorenson (2nd from bottom) in the small rack west of 1x2 was repurposed to 12V 600 mA, and was run to a terminal block on the north side of 1X1. Cables were routed underneath 1X1 and 1X2 to the terminal blocks. 12V was then routed to the PSL table and banana clip terminals were added.

  15157   Sun Jan 26 14:40:55 2020 gautamUpdateALSALS OOL noise

In preparation for resuming IFO locking activities, I measured the ALS noise with the arm lengths locked to IR, AUX laser frequencies locked to the arm lengths. Looks promising (y-axis units are Hz/rtHz).

Attachment 1: ALSnoise_20200126.pdf
ALSnoise_20200126.pdf
  15156   Sun Jan 26 13:47:00 2020 gautamUpdatePSLPMC servo characterization

Summary:

  1. I investigated the stage-by-stage transfer functions of the PMC servo up till the HV stage. See Attachment #1. There were no unexpected features.
  2. I replaced the AD602 used to implement the VGA capability. After the replacement, the gain of the VGA stage had the desired performance, see Attachment #2, Attachment #3.
  3. The servo board was re-installed and the OLTF of the PMC loop was measured. See Attachment #4.

​To avoid driving the PA85 without the HV rails connected, I removed R23. This was re-installed after my characterization.

Input stage:

Since we do the demodulation of the PMC PDH signal off this servo board, the I/F mixer output is connected to the "FP1test" front panel LEMO input.

  • A DG190 is used to enable/disable this path.
  • Initially I tried checking the enable/disable functionality by measuring the resistance across the IC's I/O pins. However, this method does not work - the resistance read off from a DMM varied from ~23 ohms in the "ON" state to ~123 ohms in the "OFF" state. While the former value is consistent with the spec, the latter is confusing.
  • But I confirmed that the switch does indeed isolate the input in the "OFF" state by injecting a signal with a function generator (100 Hz sine wave, 100mVpp) and monitoring the output on an oscilloscope.

Electronic TFs:

Using some Pomona mini-grabbers, I measured the electronic TFs between various points on the circuit. There were no unexpected features, the TFs all have the expected shape as per the annotations on the DCC schematic. I did not measure down to 0.1 Hz to confirm the low frequency pole implemented by U6, and I also didn't measure the RF low pass filter at the input stage (expected corner frequency is 1 MHz). 

VGA characterization:

After replacing the IC, I measured the transfer function between TP1 and TP2 for various values of the control voltage applied to pin 4A on the P1 connector, varying between +/- 5 V DC. 

  • Pin 9A on the P1 connector has to be grounded for the signal to be allowed to pass through the VGA. 
  • Note that there is an overall gain of -1/10 applied to the control voltage between pin 4A and pin #1 of the AD602, which is what actually sets the gain.
  • Furthermore, the input impedance of the AD602 is spec-ed to be 100 ohms. Because of the series resistance of 500 ohms from TP1 to the input of the AD602 (so that the upstream OP27 isn't overdrawn for current), the relation between the control voltage applied to Pin 4A and gain (measured between TP1 and TP2) is modified to G [dB] = 32*(-0.1 * V_pin4A) - 6. 
  • The gain behavior after the IC swap is as expected, both in terms of absolute gain, and the linearity w.r.t. the control voltage.
  • Note that in Attachment #2, each color corresponds to a different control voltage to the AD602, varying from -5V DC to +5V DC in 1V steps. 

PZT Capacitance measurement

I confirmed that the PZT capacitance is 225 nF. The measurement was made using an LCR meter connected to the BNC cable delivering the HV to the PZT, at the 1X1 rack end.

OLTF measurement

After re-soldering R23, I put the board back into its Eurocrate, and was able to lock the PMC. For subsequent measurements, the PSL shutter was closed.

  • I measured the OLTF using the usual IN1/IN2 prescription, implemented with the help of an SR560.
  • At the original PMC Servo gain of +12dB, I found that the feature at ~8kHz results in an OLTF with multiple unity gain crossings.
  • So I lowered it to +9dB. This yields an OLTF with ~60deg phase margin, ~2.3 kHz UGF. 
  • The feature that sets the gain margin is actually not any of the peaks fit by LISO, but is one of the high frequency features at ~40 kHz. At the new setting of +9dB gain, the gain margin is ~10 dB.
  • The measured TF (dots in Attachment #5) was fit with LISO (solid lines in Attachment #5) to allow inferring the out-of-loop servo noise by monitoring the in-loop noise (that plot to follow).
Attachment 1: elecTFs.pdf
elecTFs.pdf
Attachment 2: VGAchar_postFix.pdf
VGAchar_postFix.pdf
Attachment 3: VGAlinearity_postFix.pdf
VGAlinearity_postFix.pdf
Attachment 4: newOLTFs.pdf
newOLTFs.pdf
  15155   Sun Jan 26 13:30:19 2020 gautamUpdateSUSAll watchdogs tripped, now restored

Looks like a M=4.6 earthquate in Barstow,CA tripped all the suspensions. ITMX got stuck. I restored the local damping on all the suspensions just now, and freed ITMX. Looks like all the suspensions damp okay, so I think we didn't suffer any lasting damage. IMC was re-aligned and is now locked.

Attachment 1: EQ_Jan25.pdf
EQ_Jan25.pdf
  15154   Sat Jan 25 11:54:42 2020 YehonathanUpdatePSLRingdown measurements

Zero order beam PMC ringdown

On Wednesday I installed 3 PDs (see attached photos) measuring: 

1. The input light to the PMC. Flip-mirror was installed (sorry Shruti) on the beam path to the fiber coupler.

2. Reflected light from the PMC.

3. PMC transmitted light.

I connected the three PDs to the oscilloscope and the AOM driver to a function generator. I drive the AOM with a square wave going from 1V to 0V.

I slowly increased the square wave frequency. The PMC servo doesn't seem to care. I reach 100KHz - it seems excessive but still works. In any case, I get the same results doing a single shut-down from a DC level.

I download the traces. I normalize the traces but I don't rescale them (Attachment 4) so that the small extinction can be investigated.

I notice now that the PDs show the same extinction. It probably means I should have taken dark currents data for the PDs.

Also, I forgot to take the reflected data when the PMC is out of resonance with the laser which could have helped us determine the PMC transmission.

Again, the shutdown is not as sharp as I want. There is a noticeable smoothening in the transition around t = 0 which makes the fit to an exponential difficult. I suspect that the function generator is the limiting device now. I hooked up the function generator to the oscilloscope which showed similar distortion (didn't save the trace)

I try to fit the transmission PD trace to a double exponential and to Zucker model (Attachment 5).

The two exponentials model, being much less restrictive, gives a better fit but the best fit gives two identical time constants of 92ns.

The Zucker model gives a time constant of 88ns. Both of these results are consistent with more or less with the linewidth measurement but this measurement is still ridden with systematics which hopefully will become minimized IMC ringdowns.

Attachment 1: Input_beam_path.jpg
Input_beam_path.jpg
Attachment 2: Reflected_Beam_Path.jpg
Reflected_Beam_Path.jpg
Attachment 3: Transmitted_Beam_Path.jpg
Transmitted_Beam_Path.jpg
Attachment 4: PMCRingdownNormalizedRawdata.pdf
PMCRingdownNormalizedRawdata.pdf
Attachment 5: TransPDFits.pdf
TransPDFits.pdf
  15153   Fri Jan 24 17:14:01 2020 gautamUpdateALSGain blocks installed

Jordan will write up the detailed elog but in summary,

  1. Former +24V Sorensen in the AUX OMC power rack (south of 1X2) has been reconfigured to +12V DC.
  2. The voltage was routed to a bank of fusable terminal blocks on the NW corner of 1X1.
  3. An unused cable running to the PSL table was hijacked for this purpose.
  4. The ZHL-1010+ were installed on the upper shelf of the PSL table, the two gain blocks draw a total of ~600mA of current when powered.
Quote:
 

I will install these at the next opportunity, so that we can get rid of the many attenuators in this path (the main difficulty will be sourcing the required +12V DC for operation, we only have +15V available near the PSL table).

  15152   Fri Jan 24 15:42:08 2020 gautamUpdatePSLPMC servo restored

The PMC servo was re-installed at ~345pm. HV supplies were re-energized to their nominal values. I will update the results of the investigation shortly. The new nominal PMC servo gain is +9dB.

Quote:

While I have the board out, I'll try and do a thorough investigation of TFs and noise of the various stages. There is no light into the IFO until this is done.

I pulled the board out at 345pm after dialling down all the HV supplies in 1X1. I will reinstall it after running some tests.

  15151   Fri Jan 24 13:56:21 2020 JonUpdateBHDBHD optics specifications

I've started a spreadsheet for the BHD optics specifications and populated it with my best initial guesses. There are a few open questions we still need to resolve, mostly related to mode-matching:

  • PR2 replacement: What transmission do we need for a ~100 mW pickoff? Also, do we want to keep the current curvature of -700 m?
  • LO mode-matching telescope: What are the curvatures of the two mirrors?
  • Lenses: We have six of them in the current layout. What FLs do we need?

The spreadsheet is editable by anyone. If you can contribute any information, please do!

  15150   Thu Jan 23 23:07:04 2020 JonConfigurationPSLc1psl breakout board wiring

To facilitate wiring the c1psl chassis and scripting loopback tests, I've compiled a distilled spreadsheet with the Acromag-to-breakout board wiring, broken down by connector. This information is extractable from the master spreadsheet, but not easily. There were also a few apparent typos which are fixed here.

The wiring assignments at the time of writing are attached below. Here is the link to the latest spreadsheet.

Attachment 1: c1psl_feedthrough_wiring.pdf
c1psl_feedthrough_wiring.pdf c1psl_feedthrough_wiring.pdf c1psl_feedthrough_wiring.pdf c1psl_feedthrough_wiring.pdf
  15149   Thu Jan 23 22:10:01 2020 gautamUpdatePSLPMC servo pulled out

While I have the board out, I'll try and do a thorough investigation of TFs and noise of the various stages. There is no light into the IFO until this is done.

I pulled the board out at 345pm after dialling down all the HV supplies in 1X1. I will reinstall it after running some tests.

  15148   Thu Jan 23 20:08:49 2020 shrutiUpdateGeneralPLL / PM measurement of Xend NPRO PZT

Setup Update:

- No more SR 560, upgraded to LB1005 P-I controller.  Because: Elog 14687. Schematic of new setup shown in Attachment 1.

- For this, the Marconi was moved to the other (east) side of the PSL table and a power supply was also placed in the enclosure.

I think that the RF power at the mixer in this new configuration is 0 dBm (since the spectrum analyzer read ~ -20 dBm)

Progress Today:

- Turned up the HEPA to 100%, closed the PSL shutter, misaligned the ITMX, connected the LB1005 to the PZT. [The PZT has been reconnected to the X arm PDH servo, HEPA back to 20-30%]

- Tried to look for the PSL+X beat, but it was not there. Gautam identified the flipmount in the path which sorted it out (eventually), but there was no elog about itsurprise.

- After much trial, the loop seemed to lock with PI corner 1 kHz, gain ~2.9 (as read on knob), LFGL set to 90 dB. The beat note looked quite stable on the oscilloscope, but the error signal had an rms of ~100 mV (Rana pointed out that it could be the laser noise) and the lock lasted for ~1 min each time.

The parameters were similar to that in elog 14687. Why do we require such a high PI corner frequency and LFGL?

Attachment 1: Image-1.jpg
Image-1.jpg
  15147   Thu Jan 23 18:52:31 2020 gautamUpdatePSLPMC RFPD characterization

Summary:

The RF transimpedance of the PMC PDH RFPD was measured, and found to be 1.03 kV/A

Details:

With the new fiber coupled PDFR system, it was very easy to measure the response of this PD in-situ 🎉 . The usual transfer function measurement scheme was used, with the AG4395 RF out modulating the pump current of the diode laser, and the measured transfer function being the ratio of the response of the test PD to the reference PD.

I assume that the amount of light incident on the reference NF1611 photodiode and the test photodiode were equal - I don't know what the DC transimpedance of the PMC REFL photodiode is (can't find a schematic), but the DC voltage at the DC monitor point was 16.4 mV (c.f. -2.04 V for the NF1611). The assumption shouldn't be too crazy because assuming the reference PD has an RF transimpedance of 700 V/A (flat in the frequency range scanned), we get a reasonable shape for the PMC REFL photodiode's transimpedance.

The fitted parameters are overlaid in Attachment #1. The 2f notch is slightly mistuned it would appear, the ratio of transimpedance at f1/2*f1 is only ~10. The source files have been uploaded to the wiki.

Knowing this, the measured PDH discriminant of 0.064 GV/m is quite reasonable:

  • expected optical gain from modulation depth assuming a critically coupled cavity is 0.089 GW/m.
  • Assume 0.7 A/W responsivity for InGaAs.
  • Account for the fact that only 0.8 % of the reflected light reaches the PMC photodiode because of the pickoff window.
  • Account for a conversion loss of 4.5 dB in the mixer.
  • Account for the voltage division by a factor of 2 at the output of the BLP-1.9 filter due to the parallel 50 ohm termination.
  • Then, the expected PDH discriminant is 0.089e9 W/m * 0.7 A/W * 0.8e-2 * 1.03kV/A * 10^(-4.5/20) * 0.5 ~ 0.15 GV/m. This is now within a factor of ~2 of the measured value, and I assume the total errors in all the above assumed parameters (plus the cable transmission loss from the photodiode to the 1X1 rack) can easily add up to this. 

So why is this value so different from what Koji measured in 2015? Because the monitor point is different. I am monitoring the discriminant immediately after the mixer, whereas Koji was using the front panel monitor. The latter already amplifies the signal by a factor of x101 (see U2 in schematic). 

Conclusion:

I still haven't found anything that is obviously wrong in this system (apart from the slight nonlinearity in the VGA stage gain steps), which would explain why the PMC servo gain has to be lower now than 2018 in order to realize the same loop UGF.

So the next step is to characterize the RF transimpedance of the PMC RFPD.

Attachment 1: PDresp.pdf
PDresp.pdf
  15146   Thu Jan 23 16:37:14 2020 ranaUpdatePSLPMC VGA chip damaged?

doesn't seem so anomolous to me; we're getting ~25 dB of gain range and the ideal range would be 40 dB. My guess is that even thought this is not perfect, the real problem is elsewhere.

  15145   Thu Jan 23 15:32:42 2020 gautamConfigurationComputersMegatron: starts up grade

The burt snapshotting is still not so reliable - for whatever reason, the number of snapshot files that actually get written looks random. For example, the 14:19 backup today got all the snaps, but 15:19 did not. There are no obvious red flags in either the cron job logs or the autoburt log files. I also don't see any clues when I run the script in a shell. It'll be good if someone can take a look at this.

  15144   Thu Jan 23 14:37:05 2020 gautamUpdatePSLPMC VGA chip damaged?

[jordan, gautam]

Summary:

The AD602 chip which implements the overall servo gain for the PMC seems to be damaged. We should switch this out at the next opportunity.

Details:

  1. According to the PSL cross connect wiring diagram, the VME DAC that provides the control voltage to the VGA stage goes to pins 7/8 of cross connect J16.
    • Jordan and I verified that the voltage at this point [Vout], is related to the PMC_GAIN EPICS slider [dB] value according to the following relationship: V_{\mathrm{out}} = (10-\mathrm{dB})/2.
  2. On the PMC servo board, this voltage is scaled by a factor of -1/10. 
    • This was confirmed by peeking at this voltage using a DMM (I clipped onto R31) while the gain slider was varied.
    • This corresponds to +/- 1000 mV reaching the AD602.
    • However, the AD602 is rated to work with a control voltage varying between +/- 625 mV.
    • What this means is that the EPICS slider value is not the gain of the AD602 stage. The latter is given by the relation G [\mathrm{dB}] = 32 \times V_{\mathrm{G}} + 10.
    • @team PSL upgrade: this should be fixed in the database file for the new c1psl machine.
  3. Using TP1 and TP2 connected to the SR785, I measured the transfer function of the AD602 for various values of the EPICS slider.
    • Result is shown in Attachment #1.
    • I did this measurement with the PMC locked, so I'm using the in-loop error signal to infer the gain of the VGA stage.
    • As expected, the absolute value of the gain does not match that of the EPICS slider (note that the AD602 has an input impedance of 100 ohms. So the 499 ohm series resistor between TP1 and the input of AD602 makes a 1/5 voltage divider, so the gain seen between TP1 and TP2 has this factor folded in).
    • Moreover, the relative scaling of the gain for various slider values also doesn't appear to be liner.
    • For the highest gain setting of +15 dB, the servo began oscillating, so I think the apparent non-flatness of the gain as a function of frequency is an artefact of the measurement.
    • Nevertheless, my conclusion is that the IC should be changed.

I will pull the board and effect the change later today.

I pulled the board out at 345pm after dialling down all the HV supplies in 1X1. I will reinstall it after running some tests.

Attachment 1: VGAchar.pdf
VGAchar.pdf
ELOG V3.1.3-